OpenStudy (krisssssssssssss):

In studying the gene pool of a population, you find that 15% of the population is homozygous recessive (tt) for a trait. How would you use this information to solve for the frequency of the dominant T allele? a. Square root of 0.15 b. 1 – Square root of 0.15 c. 15 ÷ 100 d. 0.152

OpenStudy (krisssssssssssss):

@Jadeishere

OpenStudy (roku):

eh?

OpenStudy (sweetburger):

The Hardy Weinberg Equilibrium is represented by \[p+q=1\] and \[p^2+2pq+q^2=1\] The variables are represented by p = frequency of the dominant allele in the population q = frequency of the recessive allele in the population p2 = percentage of homozygous dominant individuals q2 = percentage of homozygous recessive individuals 2pq = percentage of heterozygous individuals. In this question we want to determine what p is.

OpenStudy (sweetburger):

We are given that \[q^2=.15\] therefore to find q we must take the square root of both sides \[q \approx .38\]. To find p we can use the equation \[p+q=1,q=.38\]. Therefore we can say \[p+.38=1,p=.62\]. What answer would correspond to the process I showed here?

alones (alones):

^ that explains everything